You are offered two different jobs selling dental supplies. One has an annual salary of $20,000 plus an end of year bonus of 1% of your total sales. The other has a salary of $15,000 and a year -end bonus of 2% of your sales. How much would you have to sell to have the jobs pay the exact same amount?
Equations: x represents:
y represents:
pay job 1 y =
pay job 2 y =
Solve the system:
Solution: If you would sell $_______________________ of equipment, both jobs would pay you $_____________________.
Analysis: If you were to sell $300,000 worth of equipment, which job would pay better? _______________. If you were to sell $700,000 worth of equipment, which job would pay better?___________________. If you sell $__________________ or more Job 2 would pay better.

Answer by mananth(16946)

Let the sales be $x
Pay job 1 --20000 +0.01x
Pay job 2-- 15000+0.02x
earnings to be equal
20000+0.01x=15000+0.02x
20000-15000=0.02x-0.01x
5000= 0.01x
/0.01
x=500,000
Sales should be $ 500,000 for eranings to be the same
20000+0.01x=15000+0.02x
plug 300000 for x
20000+300000*0.01 ---------15000+0.02*300000
23000-----------21000
Pay job I will earn more
You can work out for $700,000